2. Let Yı, ..., Yn be a random sample from an Exponential distribution with density function e-, y > 0. Let Y(1) minimum(Yi, , Yn). (a) Find the CDE of Y) b) Find the PDF of Y (c) Is θ-Yu) is an unbiased estimator of θ? Show your work. (d) what modification can be made to θ so its unbiased? Explain.
Answer #1

a) The cumulative distribution of Y_{(1)} is given by

F_{Y_{(1)}}(y_1)=P(Y_{(1)}leq y_1) = 1 - P(Y_{(1)} > y_1)

Rightarrow F_{Y_{(1)}}(y_1)=1 - P(Y_{(1)} > y_1)

Rightarrow F_{Y_{(1)}}(y_1)=1 - P(Y_1 > y_1,Y_2>y_1,....,Y_n > y_1)

Rightarrow F_{Y_{(1)}}(y_1)=1 - [P(Y_1 > y_1)]^n

Rightarrow F_{Y_{(1)}}(y_1)=1 - [1 - P(Y_1 leq y_1)]^n

Rightarrow F_{Y_{(1)}}(y_1)=1 - [1 - (1-e^{-rac{y_1}{ heta }})]^n

Rightarrow F_{Y_{(1)}}(y_1)=1 - e^{-rac{ny_1}{ heta }}, y_1 geq 0

b) The pdf of Y_{(1)} is given by

i) ду

Rightarrow f_{Y_{(1)}}(y_1) =rac{n}{ heta }e^{-rac{ny_1}{ heta }}, y_1geq 0

c) Tu u

ze dz n Jo

Rightarrow E(Y_{{(1)}})=rac{ heta }{n}*Gamma (2)

Rightarrow E(Y_{{(1)}})=rac{ heta }{n}

So, it is an unbiased estimator,

But, E(nY_{{(1)}})= heta , so nY_{{(1)}} is an unbiased estimator of heta